eBook Holt Enterprises recently paid a dividend, Do, of $3.75. It expects to have nonconstant growth of 23% for 2 years followed by a constant rate of 6% thereafter. The firms required return is 14%. a. How far away is the horizon date? is the present value of all future expected dividends at time zero. II. The terminal, or horizon, date is the date when the growth rate becomes Ш. The terminal, or horizon, date is the date when the growth rate becomes IV. The terminal, or horizon, date is the date when the growth rate becomes V. The terminal, or horizon, date is infinity since common stocks do not have nonconstant. This occurs at time zero. constant. This occurs at the beginning of Year 2 constant. This occurs at the end of Year 2. a maturity date. IV b. What is the firms horizon, or continuing, value? Do not round intermediate calculations. Round your answer to the nearest cent. c. What is the firms intrinsic value today, Po? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

a) Correct option is :

IV

I is incorrect because there are two years of non constant growth and horizon value is the start of constant growth phase.

II is incorrect because growth rate becomes constant at end of year 2

iii is incorrect because it is end of year 2 and not beginning

V is incorrect because horizon value is the start of infinite period of constant growth

b) Horinzon value = D1/(r-g) where D1 is the next period dividend:

D1=3.75*((1+0.23)^2)*1.06= $6.0137775

Value=6.0137775/(0.14-0.06)=$75.17

c) For intrinsic value we need to discount the horizon value and other dividends back to present:

Year Div Terminal value Total Discount factor Present Value
1 $       4.6125 $            4.61 0.877192982 $                 4.05
2 $       5.6734 $            75.1722 $         80.85 0.769467528 $               62.21
3 $       6.0138 $                      -  
Total $          66.2541

Stock intrinsic value=$66.25

For the perpetual growth period, terminal value of dividends = Dividend next period/(r-g) where r is the required return and g is the growth rate.

Discount factors are calculated as:1/(1+r)^n

Where r is required return and n is the year